A Church Songbook  Table Of Contents(1)  1 Jan.’94

— Contents —

1. What A Mighty God We Serve, 1

2. Finally, My Brethren, 2

3. The Lord Of Covenant, 3

4. We Will Rejoice, 4

5. Praise Him In His Sanctuary, 5

6. Rise Up, 6

7. We Are Marching In Messiah’s Band, 7

8. A Song Of Love, 8

9. Hail To The King, 9

10. The Horse And Rider, 10

11. Nothing But The Blood, 11

12. Far Above All Rule And Authority, 12

13. Sing Hallelujah, 13

14. How Lovely Is Your Dwelling Place, 14

15. Zephaniah 3:17, 15

16. Come And Let Us Go, 16

17. Behold, I Am The Lord, 17

18. Stand Up And Bless, 18

19. Jesus Is The Lord, 19

20. I Will Sing Praises, 20

21. Let God Be Magnified, 21

22. There’s A Light Shining Forth, 22

23. How Great Is Our God, 23

24. Lift Jesus Higher, 24

25. Garment Of Praise, 25

26. I Was Glad, 26

27. Oh, The Blood, 27

Property of A Church  Last Revised: January 23, 1994  Material used by permission.